OpenStudy (anonymous):

The amount of sales tax varies directly with the cost of the purchase. If the sales tax is $2.40 on a purchase of $40.00, what would be the sales tax on a purchase of $60.50?

OpenStudy (calculusxy):

You can basically use a proportion: \[\frac{ tax }{ cost }=\frac{ tax }{ cost }\]

OpenStudy (calculusxy):

Substitute the tax and the cost with the information that you already know: \[\frac{ $2.40 }{ $40.00 } = \frac{ tax }{ $60.50 }\]

OpenStudy (calculusxy):

Think of tax as "x" and you are solving for that "x."

OpenStudy (anonymous):

so think of tax as X?

OpenStudy (calculusxy):

I will rewrite the proportion: \[\frac{ $2.40 }{ $40.00 }=\frac{ x }{ $60.50 }\]

OpenStudy (calculusxy):

It's the same thing, except I replaced "tax" with "x" so that I don't confuse you.

OpenStudy (anonymous):

so x= 3.563

OpenStudy (calculusxy):

Well, I found that x=3.63

OpenStudy (calculusxy):

I think that you can recheck your work and see who is correct :)

OpenStudy (calculusxy):

Anyway, hopefully I helped you :)

OpenStudy (calculusxy):

...and you understood.